OpenStudy (anonymous):

Ellipse, center at (0, -1), length of latus rectum is 16/3, distance between foci is 4sqrt5. find the equation.

OpenStudy (anonymous):

 you have formula for your problem PF1 +PF2 = 2a and F1F2 is distance between foci, PF1 is half of latus rectum. use Pythagorean to find out PF2 from then you get a. a is the distance between center and vertices. when you got a, and have c ( foci) you have formula to figure out b by c^2 = a^2 -b^2 . then replace a, b into formula of ellipse.